OpenStudy (anonymous):

a certain mountain has an elevation of 19,380 feet. In 1923 the glacier on the peak covered 10 acres. By 2003 the glacier had melted to only 1 acre (a) assume that this glacier melted at a constant rate each year. Find the yearly rate (b) use your answer from part(a) to write a linear equation that gives the acreage A of this glacier t years past 1923. the yearly rate of change is ? acres/years what is the equation that gives the acreage of the glacier t years after 1923? A=?

OpenStudy (akashdeepdeb):

So the time passed is 80 yrs right? And the amount of ice melted = 9 acres of ice [HECK THAT'S A LOT] So rate is anything by TIME!! So the constant rate of melting = 9/80! Getting me? :) It is just like speed is rate of dstance = d/t So Ice melting rate = amt. of ice melting/years = 9/80!

OpenStudy (anonymous):

@AkashdeepDeb i came up with o.11 but im stuck after that

OpenStudy (akashdeepdeb):

So now it is kinda like distance is average Then time is TIME :P And the rate 0.11 is the speed? So what is distance? --> speed/time right? So A = 9/80/t = 9/80t :) Getting me? So that would mean that after 1 year the amount of ice that would have meleted would be? 9/80 itself! Getting me? :D

OpenStudy (anonymous):

no im totally lost @AkashdeepDeb

OpenStudy (akashdeepdeb):

Okay 9/80 means that in 80 years 9 acres were lost!

OpenStudy (akashdeepdeb):

MY BAD I AM SRY! DISTANCE = SPEED * TIME!! Haha...silly me! So Average = 9t/80!!!

OpenStudy (anonymous):

i get that much @AkashdeepDeb but i cant the (b) part of the problem

OpenStudy (akashdeepdeb):

See b)actually tells us to Find a GENERAL form! We know that in 80 years --> 9 acres are lost and in 1 year 9/80 acres are lost! Did you get till this much!?

OpenStudy (anonymous):

yes @AkashdeepDeb

OpenStudy (akashdeepdeb):

Okay so do you know that distance is given by speed * time! RIGHT?

OpenStudy (anonymous):

what is the equation that gives the acreage of the glacier t years after 1923? A=?

OpenStudy (anonymous):

o.11 right

OpenStudy (akashdeepdeb):

So this is KINDA ike that! They want us to finnd the average!! A = rate * time = 9/80 * T years so A = 9*T/80!! :D

OpenStudy (anonymous):

its a 2 answer problem

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

I get it! And the average = 9t/80 tht is the equation right????

OpenStudy (anonymous):

ok i understand now lol sry bout tht

OpenStudy (anonymous):

right @AkashdeepDeb

OpenStudy (anonymous):

yes @AkashdeepDeb

OpenStudy (akashdeepdeb):

Did you get the concept right?

OpenStudy (anonymous):

is it 0.01 @AkashdeepDeb

OpenStudy (akashdeepdeb):

It is an equation! what is 0.01 ?? :P

OpenStudy (anonymous):

what i done was 9 *o.11 then divided that by 80

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

T is a VARIABLE that means it varies always T can be 1000 , 100 , 1 , 90 etc etc. So finally you CAN right it as A = 0.11*T Got it my friend? :)

OpenStudy (anonymous):

yes